A differential wheel and axle has axles 25 cm and 10 cm in diameter. The load is 750 N and the effort required at an efficiency of 42 percent is 70 N. Calculate the wheel and diameter.



1
Expert's answer
2021-05-31T08:45:42-0400


d1=25 cm

d2 = 10 cm

Load W = 750 N

Efficiency η = 42 % = 0.42

Effort P=70 N

Velocity ratio VR "= \frac{Distance \;moved \;by \;effort}{Distance \;moved\; by \;load}"

"VR = \frac{2D}{d_1-d_2} \\

= \frac{2D}{25-10} \\

= \frac{2D}{15}=0.1333D"

Mechanical advantage "MA= \frac{Load}{Effort}"

"MA = \frac{W}{P} \\

= \frac{750}{70} = 10.714"

Efficiency "η = \frac{MA}{VR}"

"0.42 = \frac{10.714}{0.1333D} \\

0.42 \times 0.1333D = 10.714 \\

D = \frac{10.714}{0.055986}=191.37 \;cm"

Answer: 191.37 cm